Circular No. 8471

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005O IN NGC 3340 Y.-T. Chen, Graduate Institute of Astronomy, National Central Uni-

versity, Chung-Li, reports the discovery by the Taiwan Supernova Survey of an apparent supernova (mag ∼ 16.1) on unfiltered CCD images taken on Jan. 16.780 UT with the 1-m Lulin Observatory telescope. The new object had brightened to mag ∼ 15.2 on a confirming image taken on Jan. 20.650. SN 2005O is located at α = 10^{h}42^{m}18^{s}.32, δ = −0^{o}22′41′′.2 (equinox 2000.0), which is 5′′.3 east and 4′′.8 south of the center of NGC 3340. Nothing was visible at this on an image taken on 2004 Dec. 27.855 (limiting mag ∼ 18.6).

COMETS C/2004 T8, Y5, Y6, Y7, Y8 (SOHO) Further to IAUC 8469, following are the initial observations of addi-

tional Kreutz sungrazing comets found on SOHO website images.

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 T8 Oct. 4.179 12^{h}26.^{m}0 − 5^{o}30′ C3/2 HS 2005-B39
C/2004 Y5 Dec. 25.821 18 26.3 −25 53 C3/2 BZ 2005-B39
C/2004 Y6  25.983 18 25.7 −24 59 C2  HS 2005-B39
C/2004 Y7  26.267 18 27.3 −24 48 C2  HS 2005-B39
C/2004 Y8  26.738 18 33.3 −26 04 C3  HO 2005-B39

SATELLITES OF SATURN The IAU Working Group on Planetary Science Nomenclature has ap-

proved the following new designations and names of satellites of Saturn:

Saturn XXXI  Narvi  = S/2003 S 1
Saturn XXXII  Methone  = S/2004 S 1
Saturn XXXIII  Pallene  = S/2004 S 2
Saturn XXXIV  Polydeuces = S/2004 S 5

The WGPSN also notes that the spelling of three of the names of Saturn’s satellites as given on IAUC 8177 was subsequently revised. The correct spelling is as follows:

Saturn XXIII  Suttungr = S/2000 S 12
Saturn XXVII Skathi = S/2000 S 8
Saturn XXX  Thrymr = S/2000 S 7

2005 January 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green